OpenStudy (anonymous):

8x + 13y = 2 y = x- 2 Solve the system of equations A) (4/3, -2/3) B) (4/21, -38/21) C) no solution

OpenStudy (radar):

8x + 13y = 2 reaarrange y = x - 2 to read x - y =2 by legal algebra moves. x - y =2 now note that 8x + 13y = x - y or again using algebra rules 7x + 14y = 0 divide thru by 7 getting x + 2y = 0 x = -2y substitute for x in 2nd original equation getting y = -2y - 2 3y = -2 y = -2/3 Now solve for x -2/3 = x - 2 -2/3 = x -2 x = -2/3 + 2 x = 4/3

OpenStudy (radar):

did you follow with understanding?